# D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video

### v1

@@ 3.2.1 Temporal Point Probability Field
The probability field $\mathcal{P}$ models the likelihood for spatial point sampling by tracking the estimated local scene opacity $\sigma\in[0,1]$ over the course of optimization. To facilitate accurate and efficient sampling over a dynamic sequence, we divide our probability field $\mathcal{P}=(\mathcal{P}_{s},\mathcal{P}_{d})$ into two separate instances for static and dynamic regions, respectively. While the static 3D field $\mathcal{P}_{s}$ continues to track local opacity $\sigma_{s}$ assumed to be consistent over all points in time, the 4D dynamic field $\mathcal{P}_{d}$ stores additional values $\sigma_{d}(t)$ as a function of time. To enable a clean separation of foreground and background, $\mathcal{P}_{d}$ not only models the expected local opacity, but additionally keeps track of the estimated local proportion $\beta(t)$ between static and dynamic substance, as explained later on. In conjunction, our probability fields can sample explicit, time-aware point positions, while pre-marking individual points as either dynamic or static to reduce compute complexity. Given the natural imprecisions and relatively low resolution of monocular reconstruction, we find that the octree-based data structure suggested by Hahlbohm et al. [17] does not yield significant benefits due to the absence of fine geometric details. Instead, we use a fixed-resolution sparse voxel grid representation for the static and dynamic fields.
@@
Despite using a sparse data structure, starting from a fully occupied grid (i.e., all cells are sampled equally for all timestamps) is exceedingly inefficient regarding both memory consumption and model convergence. Thus, we make use of the data-driven priors described in Sec. 3.1 to find a sparse initial sampling distribution before optimization. First, we initialize the voxel grid bounds as the minimal cuboid enclosing all SfM points after performing outlier filtering based on the 95-percentile, containing a fixed total of $128^{3}$ and $N\cdot 128^{3}$ voxels for $\mathcal{P}_{s}$ and $\mathcal{P}_{d}$ respectively. Then, for every timestamp $t_{i}\in T$, we use the corresponding training camera matrix $e_{i}$ to project a set of random points in each voxel to the image plane, and gather the minimal relative distance to the estimated monocular depth $d_{i}$, and the maximum over the binary foreground segmentation mask $\omega_{i}$. We set the initial opacity-based sampling probability $\sigma_{d}(t_{i})$ of each cell to the relative deviation from the monocular depth, using the maximum over all timestamps for the static field ($\sigma_{s}=\max(\{\sigma_{d}(t)\>|\>t\in T\})$). Additionally, we carve the temporal slices of $\mathcal{P}_{d}$ using the dynamic segmentation. By thresholding the sampling probabilities, we eliminate up to 95% of voxels, resulting in an efficiently represented distribution constituting a proper local minimum for optimization.
@@
For an arbitrary camera view $e$ and timestamp $t$, we sample a set $P\in\mathbb{R}^{k\times 3}$ of $k$ explicit point positions ($P\sim\mathcal{P}\>|\>(e,t)$) for rendering a novel view. To this end, we first take the union of all voxels in the static field and the dynamic field at time $t$, taking the maximum expected opacity of voxels existing in both models. Then, we eliminate voxels based on the camera frustum by projecting each voxel center to the image plane using $e$. To obtain the final point positions $p$, we interpret the per-voxel opacity as a discrete PDF, using discrete multinomial sampling (with replacement) to propose a list of voxels, in each of which we place a single uniformly sampled 3D position. This way, more points are sampled in areas with a high expected scene density, which can iteratively be refined during optimization. For the resulting set of point positions, we further trace a binary mask $m_{p}$ indicating if a point was sampled from static or dynamic scene regions, enabling efficient model evaluation by omitting excessive queries of the dynamic appearance model.
@@
We update the sampling distribution field by tracking the blending weights assigned to every sampled explicit point during rasterization. We first apply a constant decay factor to the values of every unique voxel proposed by the multinomial sampling, and then assign it the maximum over the current value and the blending weights of all points sampled from the specific cell. For samples from dynamic voxels of $\mathcal{P}_{d}$, we apply the same procedure to the tracked static-to-dynamic ratios $\delta$ using per-point densities described in Sec. 3.2.2. Afterwards, we prune voxels from the sampling distribution by threshold: For $\mathcal{P}_{d}$, we remove cells with a low $\delta$-value, as it either describes empty space or is well handled by the static model. We further remove voxels $\mathcal{P}_{s}$ from $\mathcal{P}_{d}$ whose expected opacity $\sigma$ is low in the static 3D field and for all remaining timestamps in the dynamic field. This way, we progressively reduce memory consumption and refine importance sampling.
@@ 3.2.2 Dynamic Appearance Field
In a conventional point cloud, every point is assigned a fixed feature vector describing its appearance. As implicit neural point clouds [17] repeatedly extract new sets of point positions, they use a fast multi-resolution hash-encoded feature grid [33] $\mathcal{A}$ to model the appearance over the entire spatial sampling distribution. To represent time-conditioned appearance for dynamic scenes, we again use a static 3D and dynamic 4D hash-encoded neural feature grid $\mathcal{A}=(\mathcal{A}_{s},\mathcal{A}_{d})$. As previously described by Wang et al. [51], using an ensemble of 3D and 4D hash-grids instead of a single 4D representation effectively improves the capabilities of the dynamic model by reducing hash collisions. To model the static and dynamic scene appearance at a spatial input position $p$ at time $t$, the appearance fields yield a density $\delta\in\mathbb{R}_{0}^{+}$ and an $n_{f}$-dimensional feature vector $f\in[0,1]^{n_{f}}$:
@@
Note that we only have to query $\mathcal{A}_{d}$ for points indicated as dynamic by the binary segmentation mask $m$, which significantly reduces compute and memory. Following the formulation of D2NeRF [56], we then calculate the combined per-point density $\delta_{c}$ as the sum of static and dynamic densities, and obtain the final feature descriptor $f_{c}$ via density-weighted interpolation:
$\delta_{c}=\begin{cases}\delta_{s}+\delta_{d}&\text{if}\quad m(p)\\ \delta_{s}&\text{else}\end{cases},\qquad f_{c}=\begin{cases}\frac{\delta_{s}f_{s}+\delta_{d}f_{d}}{\delta_{c}}&\text{if}\quad m(p)\\ f_{s}&\text{else}\end{cases}.$ (2)
@@ 3.3 Neural Point Cloud Rendering
Given the time-dependent neural point cloud for an input view consisting of point positions $p$, point opacity $\sigma$, and appearance features $f$, we implement a fast differentiable point rasterizer to generate alpha ($\alpha$), depth ($d$), and 2D image feature maps $f_{\text{img}}$. Naturally, the rasterized images still contain holes, a common characteristic of classical point rendering approaches. Thus, we use a neural renderer, implemented as a light-weight convolutional neural network with a U-Net [38] architecture to fill the remaining holes, and convert the neural feature map $f_{\text{img}}$ to a human-interpretable RGB image $c$. In practice, we apply the rasterization and rendering components to the static, dynamic, and combined point cloud features separately, resulting in individual depth $d_{s,d,c}$ and color $c_{s,d,c}$ images that can be used for supervision and regularization. As the resolution of dynamic view synthesis benchmarks is usually low, we do not use any splatting or multi-resolution inputs in the rasterizer or neural renderer, which further increases performance.
@@ 3.4 Optimization
with hyperparameters $\lambda_{\text{photo}}$, $\lambda_{\text{depth}}$, $\lambda_{\text{seg}}$ and $\lambda_{\text{dist}}$. We use the same set of hyperparameters across all our experiments. The loss $\mathcal{L}_{\text{photo}}$ enforces photometric consistency between the rendered ($c$) and GT RGB image. Here, we use a weighted combination of pixel-level Cauchy loss [6], structural dissimilarity [54], and perceptual LPIPS [62] to maximize image quality. Our depth regularization loss $\mathcal{L}_{\text{depth}}$ penalizes the difference between the rasterized dynamic depth $d_{d}$ and the aligned monocular depth prior $d_{\text{gt}}$ in the binary foreground mask $\omega$:
$\mathcal{L}_{\text{depth}}=\frac{1}{\lvert\omega\rvert}\sum_{i=0}^{\lvert\omega\rvert}\omega_{i}\lvert d_{d,i}-d_{\text{gt},i}\rvert.$ (5)
The segmentation regularizer $\mathcal{L}_{seg}$ encourages a clean separation between the static and dynamic models via skewed binary entropy [56] on the dynamic ratios $\beta$:
@@
where $k$ is the skewing bias. As our last regularizer $\mathcal{L}_{dist}$, we adapt the distortion loss of Mip-NeRF360 [4, 46] to encourage thin surfaces, as it is not feasible to reliably reconstruct the time-dependent volume of a deforming object from a single monocular observation. For a detailed description of our training and loss schedule, please see our supplemental material (Sec. A.2).
@@ 4 Experimental Evaluation
We perform quantitative (Sec. 4.1) and qualitative (Sec. 4.2) evaluation and comparisons against state-of-the-art methods on two benchmark datasets: First, we evaluate our approach on the NVIDIA dynamic scenes dataset [59] (MIT license), featuring forward-facing scenes captured by a multi-view rig, monocularized by using a single frame per time to get a smooth camera trajectory. We extract training data using the procedure of RoDynRF [27], resulting in $12$ training and test views. For comparisons against DynIBaR [26] using their train/test split containing $\sim 150$ training views per scene, please see our supplemental material. Additionally, we evaluate our approach on the iPhone dataset [15] (Apache 2.0 license) which comprises challenging sequences with complex camera and scene motion trajectories, fine texture details, and (self-)occlusions. Furthermore, we provide an ablation study assessing the influence of individual pipeline components in Sec. 4.3.
@@ 4.1 Quantitative Results
Table 1: Timings on the NVIDIA dataset. We report training time and inference frame rates on the NVIDIA dataset with an image resolution of $480{\times}270$ pixels. Values for methods marked with $\dagger$ were gathered from the literature.
Method GPU Hours [h] $\downarrow$ Frame Rate [FPS] $\uparrow$
RoDynRF [27] 28 ${<}1$
HyperNeRF [35] 64 ${<}1$
DynamicNeRF [14] 74 ${<}0.1$
NSFF [25] 223 ${<}1$
DynIBaR [26] 320 ${<}1$
4D-GS [55] 1.1 45
Ours 0.27 70
@@
T-NeRF [37] 18.33 0.436 0.511
D-NeRF [37] 21.49 0.629 0.320
HyperNeRF [35] 17.60 0.377 0.502
NR-NeRF [49] 19.69 0.510 0.434
NSFF [25] 24.33 0.751 0.257
DynamicNeRF [14] 26.10 0.837 0.142
RoDynRF [27] 25.89 0.854 0.110
TiNeuVox [11] 19.74 0.501 0.412
4D-GS [55] 22.89 0.731 0.237
@@
HyperNeRF+B+D+S [35] 16.81 0.569 0.332
Nerfies+B+D+S [34] 16.45 0.570 0.339
T-NeRF [15] 15.70 0.538 0.458
T-NeRF+B+D [15] 17.06 0.573 0.390
T-NeRF+B+D+S [15] 16.96 0.577 0.379
RoDynRF [27] 17.00 0.535 0.517
4D-GS [55] 13.87 0.459 0.407
Ours 15.79 0.585 0.354
@@
We compare the performance of several related methods in Tab. 1. For all NeRF-based approaches, we report values from the literature. To further compare against a fast forward rendering approach, we train the state-of-the-art Gaussian splatting-based 4D-GS [55] on both benchmark datasets. For both 4D-GS and our method, we measure performance on a single RTX 4090 GPU. As evident from the results, our method is significantly faster than backward rendering approaches, and even trains faster than 4D-GS while achieving similar frame rates during inference. A breakdown of timings for individual method components is provided in the supplemental material. We further compare image quality based on PSNR, SSIM [54] and LPIPS [62] quality metrics on the NVIDIA and iPhone datasets in Tab. 2. To avoid inconsistencies across the reported values, we use the official results and models kindly provided by Liu et al. [27], Li et al. [26], and Gao et al. [15] to consistently recalculate all metrics using the VGG backbone for LPIPS calculation. For the iPhone dataset, we use masked metric calculation (mPSNR, mSSIM, mLPIPS) as suggested and implemented by Gao et al. [15] to account for co-visibility. Overall, we find our method is not only the fastest in terms of reconstruction and rendering speed but also achieves competitive image quality, especially in terms of perceptual image quality metrics (LPIPS). While RoDynRF and T-NeRF achieve better per-pixel PSNR scores, we find that both methods fail to preserve high frequencies, resulting in an inferior perceived image quality, as shown in the next section. Detailed per-scene results can be found in our supplement.

### v2

@@ 3.2.1 Temporal Point Probability Field
The probability field $\mathcal{P}$ models the likelihood for spatial point sampling by tracking the estimated local scene opacity $\sigma\in[0,1]$ over the course of optimization. To facilitate accurate and efficient sampling over a dynamic sequence, we divide our probability field $\mathcal{P}=(\mathcal{P}_{s},\mathcal{P}_{d})$ into two separate instances for static and dynamic regions, respectively. While the static 3D field $\mathcal{P}_{s}$ continues to track local opacity $\sigma_{s}$ assumed to be consistent over all points in time, the 4D dynamic field $\mathcal{P}_{d}$ stores additional values $\sigma_{d}(t)$ as a function of time. To enable a clean separation of foreground and background, $\mathcal{P}_{d}$ not only models the expected local opacity, but additionally keeps track of the estimated local proportion $\beta(t)$ between static and dynamic substance, as explained later on. In conjunction, our probability fields can sample explicit, time-aware point positions, while pre-marking individual points as either dynamic or static to reduce compute complexity. Given the natural imprecisions and relatively low resolution of monocular reconstruction, we find that the octree-based data structure suggested by Hahlbohm et al. does not yield significant benefits due to the absence of fine geometric details. Instead, we use a fixed-resolution sparse voxel grid representation for the static and dynamic fields.
@@
Despite using a sparse data structure, starting from a fully occupied grid (i.e., all cells are sampled equally for all timestamps) is extremely inefficient regarding both memory consumption and model convergence. Thus, we make use of the data-driven priors described in Sec. 3.1 to find a sparse initial sampling distribution before optimization. First, we initialize the voxel grid bounds as the minimal cuboid enclosing all SfM points after performing outlier filtering based on the 95-percentile, containing a fixed total of $128^{3}$ and $N\cdot 128^{3}$ voxels for $\mathcal{P}_{s}$ and $\mathcal{P}_{d}$ respectively. Then, for every timestamp $t_{i}\in T$, we use the corresponding training camera matrix $e_{i}$ to project a set of random points in each voxel to the image plane, and gather the minimal relative distance to the estimated monocular depth $d_{i}$, and the maximum over the binary foreground segmentation mask $\omega_{i}$. For the dynamic probability field $\mathcal{P}_{d}$, we set the initial opacity-based sampling probability $\sigma_{d}(t_{i})$ of each cell based on the relative deviation ($d\in[0,1]$) from the monocular depth: $\sigma_{d}(t_{i})=(1-d(t_{i}))^{x}$ with $x$ being a free hyperparameter, to initially sample more points close to the expected surface. For the static probability field ($\mathcal{P}_{s}$), we instead use the maximum over all timestamps ($\sigma_{s}=\max\{\sigma_{d}(t_{i})\>|\>t_{i}\in T\}$).
Additionally, we carve the temporal slices of $\mathcal{P}_{d}$ using the dynamic segmentation. By thresholding the sampling probabilities, we eliminate up to 95% of voxels, resulting in an efficiently represented distribution constituting a proper local minimum for optimization.
@@
For an arbitrary camera view $e$ and timestamp $t$, we sample a set $P\in\mathbb{R}^{k\times 3}$ of $k$ explicit point positions ($P\sim\mathcal{P}\>|\>(e,t)$) for rendering a novel view. To this end, we first take the union of all voxels in the static field and the dynamic field at time $t$, taking the maximum expected opacity of voxels existing in both models. Then, we eliminate voxels based on the camera frustum by projecting each voxel center to the image plane using $e$. To obtain the final point positions $p$, we interpret the per-voxel opacity as a discrete PDF, using discrete multinomial sampling (with replacement) to propose a list of voxels, in each of which we place a single uniformly sampled 3D position. This way, dense point clouds are sampled in clusters of high expected scene density, while empty or homogeneous static areas are sampled sparsely, enabling the renderer to infer depth ordering and occlusions via alpha-blending. For the resulting set of point positions, we further trace a binary mask $m_{p}$ indicating if a point was sampled from static or dynamic scene regions, enabling efficient model evaluation by omitting excessive queries of the dynamic appearance model.
@@
We update the sampling distribution field by tracking the blending weights $(b_{i}\in\mathcal{B})$ assigned to every sampled explicit point during rasterization. We first apply a constant decay factor $(\gamma)$ to the values of every unique voxel proposed by the multinomial sampling, and then assign it the maximum over the current value and the blending weights of all points sampled from the specific cell:
$\sigma_{s}=\max\{\sigma_{s}\gamma,\ max\{\mathcal{B}\}\},$ (1)
where $\mathcal{B}$ is the set of blending weights, i.e., the visibility of all points sampled from this cell, which is calculated during alpha-blending at the rasterization step. For samples from dynamic voxels of $\mathcal{P}_{d}$, we apply the same procedure to the tracked static-to-dynamic ratios $\delta$ using per-point densities described in Sec. 3.2.2. Afterwards, we prune voxels from the sampling distribution by threshold: For $\mathcal{P}_{d}$, we remove cells with a low sampling probability, as it either describes empty space or is well handled by the static model. We further remove static voxels from $\mathcal{P}_{s}$ whose expected opacity $\sigma$ is low both in the static 3D field and all timestamps in the dynamic field. This way, we progressively reduce memory consumption and refine importance sampling.
@@ 3.2.2 Dynamic Appearance Field
In a conventional point cloud, every point is assigned a fixed feature vector describing its appearance. As implicit neural point clouds [HFK24] repeatedly extract new sets of point positions, they use a fast multi-resolution hash-encoded feature grid [MESK22] $\mathcal{A}$ to model the appearance over the entire spatial sampling distribution. To represent time-conditioned appearance for dynamic scenes, we again use a static 3D and dynamic 4D hash-encoded neural feature grid $\mathcal{A}=(\mathcal{A}_{s},\mathcal{A}_{d})$. As previously described by Wang et al. [WCW23], using an ensemble of 3D and 4D hash-grids instead of a single 4D representation effectively improves the capabilities of the dynamic model by reducing hash collisions. To model the static and dynamic scene appearance at a spatial input position $p$ at time $t$, the appearance fields yield a density $\delta\in\mathbb{R}_{0}^{+}$ and an $n_{f}$-dimensional feature vector $f\in[0,1]^{n_{f}}$:
@@
Note that we only have to query $\mathcal{A}_{d}$ for points indicated as dynamic by the binary segmentation mask $m$, which significantly reduces compute and memory. Following the formulation of D2NeRF [WZT22], we then calculate the combined per-point density $\delta_{c}$ as the sum of static and dynamic densities, and obtain the final feature descriptor $f_{c}$ via density-weighted interpolation:
$\delta_{c}=\begin{cases}\delta_{s}+\delta_{d}&\text{if}\quad m(p)\\ \delta_{s}&\text{else}\end{cases},\qquad f_{c}=\begin{cases}\frac{\delta_{s}f_{s}+\delta_{d}f_{d}}{\delta_{c}}&\text{if}\quad m(p)\\ f_{s}&\text{else}\end{cases}$ (3)
@@ 3.3 Neural Point Cloud Rendering
Given the time-dependent neural point cloud for an input view consisting of point positions $p$, point opacity $\sigma$, and appearance features $f$, we implement a fast differentiable point rasterizer to generate alpha ($\alpha$), depth ($d$), and 2D image feature maps $f_{img}$. Naturally, the rasterized images still contain holes, a common characteristic of classical point rendering approaches. Thus, we use a neural renderer, implemented as a light-weight convolutional neural network with a U-Net [RFB15] architecture. This network translates the sparse rasterized feature image ($f_{img}$) to a dense output image ($c$), thus inferring human-interpretable RBG color from the higher dimension point cloud features while filling the remaining holes based on neighborhood information. In practice, we apply the rasterization and rendering components to the static, dynamic, and combined point cloud features separately, resulting in individual depth $d_{s,d,c}$ and color $c_{s,d,c}$ images that can be used for supervision and regularization. As the resolution of dynamic view synthesis benchmarks is usually low, we do not use any splatting or multi-resolution inputs in the rasterizer or neural renderer, which further increases performance.
@@ 3.4 Optimization
with hyperparameters $\lambda_{\text{photo}}$, $\lambda_{\text{depth}}$, $\lambda_{\text{seg}}$ and $\lambda_{\text{dist}}$. We use the same set of hyperparameters across all our experiments. The loss $\mathcal{L}_{\text{photo}}$ enforces photometric consistency between the rendered ($c$) and GT RGB image. Here, we use a weighted combination of pixel-level Cauchy loss [BA96], structural dissimilarity [WBSS04], and perceptual LPIPS [ZIE18] to maximize image quality. Our depth regularization loss $\mathcal{L}_{\text{depth}}$ penalizes the difference between the rasterized dynamic depth $d_{d}$ and the aligned monocular depth prior $d_{\text{gt}}$ in the binary foreground mask $\omega$:
$\mathcal{L}_{\text{depth}}=\frac{1}{\lvert\omega\rvert}\sum_{i=0}^{\lvert\omega\rvert}\omega_{i}\lvert d_{d,i}-d_{\text{gt},i}\rvert$ (6)
The segmentation regularizer $\mathcal{L}_{seg}$ encourages a clean separation between the static and dynamic models via skewed binary entropy [WZT22] on the dynamic ratios $\beta$:
@@
where $k$ is the skewing bias. As our last regularizer $\mathcal{L}_{dist}$, we adapt the distortion loss of Mip-NeRF360 [BMV22, SSC22] to encourage thin surfaces, as it is not feasible to reliably reconstruct the time-dependent volume of a deforming object from a single monocular observation. For a detailed description of our training and loss schedule, please see our supplemental material.
Method GPU Hours [h] $\downarrow$ Frame Rate [FPS] $\uparrow$ Frame Rate [FPS] $\uparrow$
(NVIDIA dataset) ($480{\times}270$) ($1920{\times}1080$)
RoDynRF [LGM23] 28 ${<}1$
HyperNeRF [PSH21] 64 ${<}1$
DynamicNeRF [GSKH21] 74 ${<}0.1$
NSFF [LNSW21] 223 ${<}1$
DynIBaR [LWC23] 320 ${<}1$
Marbles [SHU24] 3.5 200+
CasualFVS [LZBM25] 0.25 48
4D-GS [WYF24] 1.1 45 30
Ours 0.27 78 22
$\dagger$ Values gathered from literature – Values not reported
Table 1: Performance Comparison. We report training times on the NVIDIA dataset, and inference frame rates for training- and FHD resolution.
TiNeuVox RoDynRF 4D-GS Ours Ground Truth
[…]
Figure 3: Comparison on the iPhone dataset. We compare on five challenging scenes of the iPhone dataset. Our method better preserves the object shapes and fine details despite complex deformations. Methods indicated by $\dagger$ are the highest quality versions provided by Gao et al. [GLT22]. The areas highlighted in red indicate lack of co-visibility between training and validation.
@@ 4 Experimental Evaluation
We perform quantitative (Sec. 4.1) and qualitative (Sec. 4.2) evaluation and comparisons against state-of-the-art methods on two benchmark datasets: First, we evaluate our approach on the NVIDIA dynamic scenes dataset [SYKG20], featuring forward-facing scenes captured by a multi-view rig, monocularized by using a single frame per time to get a smooth camera trajectory. We extract training data using the procedure of RoDynRF [LGM23], resulting in $12$ training and test views. Additionally, we evaluate our approach on the iPhone dataset [GLT22] which comprises challenging sequences with complex camera and scene motion trajectories, fine texture details, and (self-)occlusions. We further show qualitative results including foreground-background separation on high-resolution (1080p) in-the-wild scenes from the DAVIS dataset [PPTM16], and provide an ablation study assessing the influence of individual pipeline components in Sec. 4.3. For further experiments, including comparisons against DynIBaR [LWC23] and results on the HyperNeRF dataset [PSH21], please see our supplemental PDF.
@@ 4.1 Quantitative Results
We compare the performance of our approach to several related benchmark methods. In addition to recent NeRF-based approaches, we also compare against the official implementation of 4D-GS [WYF24], a fast state-of-the-art forward rendering approach based on Gaussian splatting. For both 4D-GS and our method, we measure frame rates using the procedure described by SMERF [DHR24], rendering the test set $100$ times using a single RTX 4090 GPU. Please note that we can not report all numbers for concurrent (unpublished at time of submission) methods [LZBM25, WYG24, SHU24, LWH24] due to missing source code or differences in evaluation procedure (e.g., training resolution or camera calibration on the iPhone dataset).
@@
T-NeRF [PCPMMN21] 18.33 0.436 0.511
D-NeRF [PCPMMN21] 21.49 0.629 0.320
HyperNeRF [PSH21] 17.60 0.377 0.502
NR-NeRF [TTG21] 19.69 0.510 0.434
NSFF [LNSW21] 24.33 0.751 0.257
DynamicNeRF [GSKH21] 26.10 0.837 0.142
RoDynRF [LGM23] 25.89 0.854 0.110
TiNeuVox [FYW22] 19.74 0.501 0.412
4D-GS [WYF24] 22.89 0.731 0.237
@@
HyperNeRF+B+D+S [PSH21] 16.81 0.569 0.332
Nerfies+B+D+S [PSB21] 16.45 0.570 0.339
T-NeRF [GLT22] 15.70 0.538 0.458
T-NeRF+B+D [GLT22] 17.06 0.573 0.390
T-NeRF+B+D+S [GLT22] 16.96 0.577 0.379
RoDynRF [LGM23] 17.00 0.535 0.517
Marbles [SHU24] 16.72 0.418
4D-GS [WYF24] 13.87 0.459 0.407
Ours 16.04 0.589 0.348
@@
As evident from Tab. 1, our method is significantly faster than backward rendering approaches, and even trains faster than 4D-GS while achieving similar frame rates during inference. A breakdown of timings for individual method components is provided in the supplemental material. We further compare image quality based on PSNR, SSIM [WBSS04], and LPIPS [ZIE18] quality metrics on the NVIDIA and iPhone datasets in Tab. 2. To avoid inconsistencies across the reported values, we use the official results and models kindly provided by Liu et al. [LGM23], Li et al. [LWC23], and Gao et al. [GLT22] to consistently recalculate all metrics, using the VGG backbone for LPIPS calculation. On the iPhone dataset, we instead use masked metric calculation (mPSNR, mSSIM, mLPIPS) as suggested and implemented by Gao et al. [GLT22] to account for co-visibility. For our method, we further report values for two versions, trained with our monocular depth alignment (Ours) and the provided ${}_{\text{Lidar}}$). Overall, we find our method is not only the fastest in terms of reconstruction and rendering speed but also achieves competitive image quality, especially in terms of perceptual image quality metrics (LPIPS). While RoDynRF and T-NeRF achieve better per-pixel PSNR scores, we find that both methods fail to preserve high frequencies, resulting in an inferior perceived image quality, as shown in the next section. Detailed per-scene results can be found in our supplement.